OpenStudy (anonymous):

How to differentiate sin(cos(x^2)) with respect to x?

OpenStudy (anonymous):

\[-2 x \sin \left(x^2\right) \cos \left(\cos \left(x^2\right)\right) \]

OpenStudy (anonymous):

How it came like that?

OpenStudy (anonymous):

Refer to the attachemnts which came from WolframAlpha.com

OpenStudy (anonymous):

attachemnts should have been typed attachments